OpenStudy (gabby1214):

solve x/6=-8 solution x=

OpenStudy (anonymous):

What do you get when you multiply both sides by 6?

OpenStudy (michele_laino):

If I multiply both sides by \(6\) I got: \[\huge \frac{x}{6} \cdot 6 = \left( { - 8} \right) \cdot 6\] please simplify

OpenStudy (gabby1214):

is it -48??